# Back from the dead: AT2019aalc as a candidate repeating TDE in an AGN

### v1

@@ 3.2 Infrared
The Palomar 200-inch (P200) near-infrared telescope equipped with the Wide Field Infrared Camera (WIRC, Wilson et al. 2003) observed AT2019aalc on 2023-07-05. The imaging was obtained in the J-, H,- and Ks-bands. The data was processed with the Swarp software (Bertin et al. 2002). We performed background subtraction using the photutils.background class and measured the flux using a circular aperture with a radius of pixels. The corresponding zero-point magnitudes for this aperture size were calculated using the automatic pipeline for each filter. The error bars are the computed Poisson and background error of the photometry222https://www.stsci.edu/hst/instrumentation/wfc3/data-analysis/photometric-calibration/ir-photometric-calibration and we also took into account the uncertainty of the derived zero-point magnitudes. Finally, we corrected for galactic dust extinction.
We obtained mid-infrared photometry from archival observations of the Wide-Field Survey Explorer (WISE, Wright et al. 2010), which is continuously monitoring the sky at (W1) and, due to the NEOWISE project (Mainzer et al. 2011), at (W2). It visits any point in the sky roughly every six months for about ten times. Using timewise (Necker & Mechbal 2024; Necker et al. 2024), we downloaded all available multi-epoch photometry data from the AllWISE data release and single-epoch photometry from the NEOWISE-R data releases and stacked the data per visit to obtain robust measurements. All selected single-exposure photometry data points are within arcsecond of the nucleus of the host. The resulting light curve spans almost years, starting in February until July .
@@ 3.3 X-ray
The Swift’s X-ray Telescope (XRT, Burrows et al. 2005) observations were obtained simultaneously with the UVOT observations. The observations were performed in photon counting (PC) mode. For each of the epochs, we performed the data reduction using the Space Science Data Center (SSDC) interactive archive in the following way. First, cleaned event-files were produced using the xrtpipeline task. The exposure map, used to correct for the loss of flux caused by some of the CCD pixels not being used to collect data, was created with the xrtexpomap tool. The ximage tool allowed us to determine the source detection level and the source position. In case of non-detections, upper limits for the count rate were derived and converted to flux using the WebPIMMS333https://heasarc.gsfc.nasa.gov/cgi-bin/Tools/w3pimms/w3pimms.pl tool and assuming the mean power-law index of the observations yielded in detections. When our target source was detected above the level, we ran the swxrtdas tool in order to extract the source and the background spectra. Source counts were retrieved from a circle with a radius of pixel ( arcseconds), while background counts were extracted using an annular region with pixel inner/outer radius. In addition, Ancillary Response Files (ARFs) i.e. effective response files for the extracted spectra and Response Matrices Files (RMF), that are used to link the instrumental channel scale with the physical wavelength scale, were automatically created. Finally, the grppha tool of the heasoft 6.32 software package was used to perform background subtraction using the source and background spectra and to create a readable spectrum for spectral fitting tools. The relevant ARF and RMF files were applied during this step. We rebinned each spectrum to have a minimum of counts per spectral bin. Channels below keV were ignored.
@@
We fitted the individual X-ray spectra using the XSPEC fitting environment (Arnaud 1996b) in order to be able to study the X-ray flux evolution of AT2019aalc. First, we have determined the needed Galactic neutral hydrogen column density, , based on the HI4PI survey (HI4PI Collaboration et al. 2016). Then, we loaded the source spectrum and converted the instrumental Channel scale to ’physical’ wavelength scale. The spectrum was then fitted using Cash-statistic (C-stat, Cash 1976). To find the best-fitting model, we first fitted the stacked spectrum with a absorbed power-law (Tbabs*powerlaw in XSPEC) model as most of the AGN are dominated by a simple power-law and the last XMM data of AT2019aalc before our monitoring program was best-fit with a power-law model (Pasham 2023). We performed renormalization and ignored bad data points. When fitting the model, the galactic column density was fixed whilst the power-law index and the normalization were free to vary. The resultant fit quality is C-Stat for degrees of freedom. We than added a blackbody model component, to take into account the soft excess. The blackbody temperature and the normalization of the new component were also free to vary and we fitted again which resulted in a better fit with C-Stat for degrees of freedom. We implemented the likelihood ratio test (lrt in XSPEC) to estimate the statistical significance of the inclusion of the blackbody component. This test runs a given number of simulations of datasets based on the two models provided (which are the simple power-law and the more complex power-law+blackbody models in our case) and calculates the likelihood ratio for the more complex model relative to the simpler one. After running simulations, the test resulted only in smaller likelihood differences than the observed one () clearly implying that the difference is unlikely to be due to random fluctuations. Therefore, accounting for the soft X-ray excess results in a significantly better fit than a normal power-law fit to model the stacked spectrum. The source luminosity of each dataset was determined with adding the clumin component in front of the model components. We fixed the normalization in the additive models (power-law and blackbody) and performed the fit again. Finally, we derived the confidence level () error of the luminosity and the power-law index and, when it was relevant, the blackbody temperature. When fitting the individual datasets, a simple power-law model was enough to adequately describe the data and the soft excess becomes significant only in the stacked spectrum. Only one of the individual datasets is better fit with a blackbody model than with a power-law+blackbody or a simple power-law models.
@@ 3.4.1 VLBI observations
We observed AT2019aalc with the European VLBI Network (EVN) and the enhanced Multi-Element Remotely Linked Interferometer Network (e-MERLIN, Garrington et al. 2004). Our very long baseline interferomety (VLBI) maps enable us to study the pc-scale radio structure of AT2019aalc. These observations were carried out at GHz (L-band) on 2023 June 09 (project code: EV027, PI: Veres). Each of the eight intermediate-frequency channels (IFs) were divided into thirty-two kHz spectral channels, resulting in a total bandwidth of MHz. The whole observation time was hr. The experiment was performed in phase-referencing mode with -min duty cycles (with spending min on the target during each cycle) which resulted in an on-source time of h including fringe-finder scans and slewing. The following antennas participated in the observations: Cm: Cambridge, UK; Da: Darnhall, UK; De: Defford, UK, Ef - Effelsberg, Germany; Hh - Hartebeesthoek, South Africa; Jb - Jodrell Bank Mk2, UK; Kn - Knockin, UK; MC - Medicina, Italy; Pi - Pickmere, UK; O8 - Onsala, Sweden; T6 - Tianma, China; Tr - Toruń, Poland; Ur - Urumqi, China; Wb - Westerbork, The Netherlands. The data were recorded in left and right circular polarizations with a data rate of Gbit s-1. The correlation was done at the EVN Data Processor (Keimpema et al. 2015) at the Joint Institute for VLBI European Research Infrastructure Consortium (Dwingeloo, The Netherlands) with an averaging time of s. The Compact Symmetric Object (CSO) 15210430 (a.k.a. 4C 04.51) at an angular distance of was observed as phase-reference calibrator. Earlier EVN observations of this calibrator source taken at GHz resolved two lobes, with a hint of a jet-like emission (Xiang et al. 2002). In addition, the following radio sources were observed as fringe-finders: J14301043, J15500527 and J17510939. The bright CSO J14072827 (a.k.a. OQ208) was observed as amplitude calibrator and in order to calibrate the Merlin antennas properly.
@@
The data were calibrated with the U.S. National Radio Astronomy Observatory Astronomical Image Processing System (AIPS, Greisen 1990) software package following the standard procedures. We performed time flagging using the .uvflg ancilliary table. The amplitude calibration has been done using the measured system temperatures444Wb, T6, Ur and the e-MERLIN (Cm, Da, De, Kn, Pi) antennas had no or bad system temperature measurements, we used nominal values for amplitude calibration. and gain curves of each telescope contained in the .antab ancilliary table. Since our observations were conducted at below Ghz, we corrected for the dispersive ionospheric delays. Then, we corrected for the parallactic angle and performed global fringe-fitting for the calibrator sources. We imaged the phase-calibrator with the Difmap software (Shepherd et al. 1994), using the hybrid mapping technique with clean deconvolution (Högbom 1974). It included several iterations of cleaning and phase self-calibration before we performed an overall amplitude self-calibration to obtain the antenna-specific gain correction factors. We then applied these amplitude corrections for telescopes with deviation from the gain solutions in AIPS. Using the clean component (CC) model we built up in Difmap, we repeated the fringe-fitting in AIPS for the phase-reference calibrator source to correct for the source structure since it contributes to the measured interferometric phases. Finally, we interpolated the solutions of the last fringe-fitting to the data of our target source. As a final step, we exported the calibrated visibility file from AIPS to Difmap for imaging purposes described in the following.
After flagging the outlier data points, we imaged the target source using the hybrid mapping technique including the clean algorithm and phase-only self-calibration. Then, we obtained the antenna-specific gain correction factors. At the end of the procedure, several steps of amplitude and phase self-calibration were performed, starting from the total observation time, gradually for shorter and shorter solution time intervals.
A simple elliptical Gaussian modelfit component (Pearson 1995) was fitted directly to the visibility data. This way we can describe the brightness distribution of the radio source for a quantitative analysis and to reveal any possible structure of it. Finally, to decrease the noise level, we performed clean iterations in the residual image with a small loop gain of .
@@ 3.4.2 ATCA monitoring
The Australia Telescope Compact Array (ATCA)555https://csiropedia.csiro.au/australia-telescope-compact-array/, located at Narrabri, New South Wales, Australia, was used to monitor AT2019aalc. The observations were conducted between 2023-06-28 and 2024-05-04, utilizing the array in the 6km (6A and 6D) and (H168) configuration. The observations were conducted at frequencies of GHz, GHz, GHz, GHz and GHz. The flux calibrator was the standard ATCA primary calibrator, 1934–638, and 1548+056 is used as phase calibrator. The observations are summarized in Table LABEL:table:atca.
The data reduction procedures were based on the Multichannel Image Reconstruction, Imaging Analysis and Display (miriad; Sault et al. 1995). The data reduction procedure was conducted in accordance with standard protocol. In order to mitigate the impact of radio frequency interference (RFI) during flux and phase calibration, the interactive flagging tool bflag was utilised. Furthermore, automated flagging routines pgflag were applied to address interference for the source. Any corrupted data identified during calibration was manually flagged.
@@
Date | Exp. time / min | Configuration | Central Freq. / GHz | Bandwidth / GHz | Flux Calibrator | Phase Calibrator |
28.06.2023 | 330 / 180 | 6D | 2.1 / 7.0 | 1 / 4 | 1934-638 | 1548+056 |
30.06.2023 | 90 / 180 | 6D | 2.1 / 7.0 | 1 / 4 | 1934-638 | 1548+056 |
08.07.2023 | 120 / 120 | 6D | 2.1 / 7.0 | 1 / 4 | 1934-638 | 1548+056 |
19.07.2023 | 210 | 6D | 2.1 | 1 | 1934-638 | 1548+056 |
09.08.2023 | 120 / 150 | 6D | 2.1 / 7.0 | 1 / 4 | 1934-638 | 1548+056 |
19.10.2023 | 180 / 180 | H168 | 2.1 / 7.0 | 1 / 4 | 1934-638 | 1548+056 |
09.02.2024 | 60 / 60 / 60 | 6A | 2.1 / 7.0 | 1 / 4 / 4 | 1934-638 | 1548+056 |
29.03.2024 | 60 / 60 / 60 | 6A | 2.1 / 7.0 / 17.0 | 1 / 4 / 4 | 1934-638 | 1548+056 |
@@ 3.5 Spectroscopic observations
The Low Resolution Imaging Spectrograph (LRIS; Oke et al. 1995) mounted on the Keck-I telescope at Maunakea measured the spectrum of AT2019aalc, on 2021-07-06 (PI: Kulkarni), almost exactly at half time between the two optical flares. At that time the source was not significantly detected with ZTF as indicated by the forced photometry results. Later, during the second optical flare, we obtained more optical spectra of AT2019aalc using the following instruments: LRIS (PI: Kulkarni), DeVeny Spectrograph mounted on the Lowell Discovery Telescope (LDT) in Arizona (PI: Hammerstein) and the Double Beam Spectrograph (DBSP, Oke & Gunn 1982) mounted on the inch telescope at Palomar Observatory, California (PIs: Kulkarni and Kasliwal). These optical spectra were reduced following the standard procedures using the automated reduction pipeline of the LRIS (Perley 2019), the software package PypeIt (Prochaska et al. 2020) for DeVeny, and the software packages DBSP DRT (Mandigo-Stoba et al. 2022) and PypeIt (Prochaska et al. 2020) for the DBSP. We normalized the flux of each spectrum and plotted it versus the rest-frame wavelength. Details of these observations of AT2019aalc are summarized in Table LABEL:table:sp. The obtained spectra cover a timescale of a year and allow us to study the spectroscopic evolution of the transient during its optical re-brightening. We describe the features of each spectra in Subsect. 4.5 and discuss the remarkable lines in Subsect. 5.1.
Date | Telescope | Instrument | Exp. | |
time | ||||
(d) | (s) | |||
Post-flare-1 | ||||
2021-07-06 | Keck-I | LRIS | 300 | |
Over-flare-2 | ||||
2023-06-15 | Keck-I | LRIS | 300 | |
2023-06-22 | LDT | DeVeny | 800 | |
2023-07-11 | Palomar Hale m | DBSP | 180 | |
2023-08-17 | Palomar Hale m | DBSP | 900 | |
2024-05-21 | Palomar Hale m | DBSP | 1200 |
@@
To determine the type of AGN, we performed a stellar population synthesis fit of the host galaxy’s SDSS spectrum with (pPXF; Cappellari & Emsellem 2004; Cappellari 2017, 2023) and the newest GALAXEV (Bruzual & Charlot 2003) models (CB2019), assuming a Kroupa (2001) initial mass function (IMF) and an upper mass limit of , as templates. Beforehand, the spectrum was corrected for Milky Way extinction using the dust reddening maps by Schlafly & Finkbeiner (2011) and Cardelli et al. (1989) extinction curve. Furthermore, it was redshift corrected and logarithmically rebinned. We included several emission lines to the fit and assumed four kinematic components: One for all stellar templates, one for all forbidden lines, one for all allowed emission lines, and a last one for a second, broad component of all Balmer lines up to . Additionally, we included two separate dust components, each one with the extinction curve of Cardelli et al. (1989): a first one for stellar, and a second one for nebular emission. We therefore fixed the Balmer line ratio and limited doublets of the preinstalled lines to their theoretical ratio using the keywords tie_balmer and limit_doublets of pPXF. However, we only fixed the ratio for the strongest lines of the Balmer series up to to avoid possible contamination of by the [Ne iii] line. For the fit with pPXF we included additive and multiplicative polynomials up to tenth order (degree=10, mdegree=10). In order to correct for intrinsic dispersion caused by the spectrograph, we assume a resolving power of at and at .666https://www.sdss4.org/dr17/spectro/spectro_basics/ Linear interpolation is used to calculate values for wavelengths in between.
The SPS fit of the latest DBSP spectrum was conducted in a manner analogous to the one applied to the SDSS. However, additional kinematic components were required to account for the internal structure in order to achieve a satisfactory result. Each emission line with an excitation potential above is treated as a separate component. The [S ii] doublet is fitted as an additional component as well.

### v2

@@ 3.2 Infrared
The Palomar 200-inch (P200) near-infrared telescope equipped with the Wide Field Infrared Camera (WIRC, Wilson et al. 2003) observed AT2019aalc on 2023-07-05. The imaging was obtained in the J-, H,- and Ks-bands. The data was processed with the Swarp software (Bertin et al. 2002). We performed background subtraction using the photutils.background class and measured the flux using a circular aperture with a radius of pixels. The corresponding zero-point magnitudes for this aperture size were calculated using the automatic pipeline for each filter. The error bars are the computed Poisson and background error of the photometry222https://www.stsci.edu/hst/instrumentation/wfc3/data-analysis/photometric-calibration/ir-photometric-calibration and we also took into account the uncertainty of the derived zero-point magnitudes.
We obtained mid-infrared photometry from archival observations of the Wide-Field Survey Explorer (WISE, Wright et al. 2010), which was continuously monitoring the sky at (W1) and, due to the NEOWISE project (Mainzer et al. 2011), at (W2). It visits any point in the sky roughly every six months for about ten times. Using timewise (Necker & Mechbal 2024; Necker et al. 2024), we downloaded all available multi-epoch photometry data from the AllWISE data release and single-epoch photometry from the NEOWISE-R data releases and stacked the data per visit to obtain robust measurements. All selected single-exposure photometry data points are within arcsecond of the nucleus of the host. The resulting light curve spans almost years, starting in February until July 2024.
@@ 3.3 X-ray
The X-ray Telescope (XRT, Burrows et al. 2005) observations were obtained simultaneously with the UVOT observations. The observations were performed in photon counting mode. We performed the data reduction using the Space Science Data Center (SSDC) interactive archive in the following way. First, cleaned event-files were produced using the xrtpipeline and xrtexpomap tasks. The ximage tool allowed us to determine the source detection level and its position. In case of non-detections, upper limits for the count rate were derived and converted to flux using WebPIMMS333https://heasarc.gsfc.nasa.gov/cgi-bin/Tools/w3pimms/w3pimms.pl and assuming the mean power-law index of the observations that resulted in detections. When our target source was detected above the level, we ran swxrtdas in order to extract the source and the background spectra. Source counts were retrieved from a circle with a radius of pixel ( arcseconds), while background counts were extracted using an annular region with pixel inner/outer radius. Finally, the grppha tool of the heasoft 6.32 software package was used to perform background subtraction using the source and background spectra. We binned each spectrum to have a minimum of counts per spectral bin.
@@
We fit the individual X-ray spectra using the XSPEC fitting environment (Arnaud 1996b). First, we have determined the Galactic neutral hydrogen column density, , based on the HI4PI survey (HI4PI Collaboration et al. 2016). The spectrum was then fitted using Cash-statistic (C-stat, Cash 1976). To find the best-fitting model, we first fitted the stacked spectrum with a absorbed power-law (Tbabs*powerlaw in XSPEC) model as the last XMM data of AT2019aalc before our monitoring program was best-fit with a power-law model (Pasham 2023). When fitting the model, the galactic column density was fixed whilst the power-law index and the normalization were free to vary. The resultant fit quality is C-Stat for degrees of freedom. We than added a blackbody model component, to take into account a possible soft excess. The blackbody temperature and the normalization of the new component were also free to vary and we fit again which resulted in a better fit with C-Stat for degrees of freedom. We implemented the likelihood ratio test (lrt in XSPEC) to estimate the statistical significance of the inclusion of the blackbody component. After running simulations, the test resulted only in smaller likelihood differences than the observed one clearly implying that the difference is unlikely to be due to random fluctuations. Therefore, accounting for the soft X-ray excess results in a significantly better fit.
@@ 3.4.1 VLBI observations
We observed AT2019aalc with the European VLBI Network (EVN) and the enhanced Multi-Element Remotely Linked Interferometer Network (e-MERLIN, Garrington et al. 2004). These observations were carried out at GHz (L-band) on 2023 June 09 (project code: EV027, PI: Veres). The observation time was hr. The experiment was performed in phase-referencing mode with -min duty cycles (with spending min on the target during each cycle) which resulted in an on-source time of h including fringe-finder scans and slewing. The data were recorded in left and right circular polarizations with a data rate of Gbit s-1. The correlation was done at the EVN Data Processor (Keimpema et al. 2015) at the Joint Institute for VLBI European Research Infrastructure Consortium (Dwingeloo, The Netherlands) with an averaging time of s. The Compact Symmetric Object (CSO) 15210430 (a.k.a. 4C 04.51) at an angular distance of was observed as phase-reference calibrator.
@@
The data were calibrated with the U.S. National Radio Astronomy Observatory Astronomical Image Processing System (AIPS, Greisen 1990) software package following the standard procedures including time flagging, amplitude calibration, correction for the dispersive ionospheric delays and global fringe-fitting of the calibrator sources. We imaged the phase-calibrator with the Difmap software (Shepherd et al. 1994), using the hybrid mapping technique (Högbom 1974). Using the clean component (CC) model we built up in Difmap, we repeated the fringe-fitting in AIPS for the phase-reference calibrator source to correct its structure. Finally, we interpolated the solutions of the last fringe-fitting to the data of our target source. We imaged the target source using the hybrid mapping technique in Difmap. At the end of the procedure, several steps of amplitude and phase self-calibration were performed. A simple elliptical Gaussian modelfit component (Pearson 1995) was fitted directly to the visibility data. Finally, to decrease the noise level, we performed clean iterations in the residual image with a small loop gain of .
@@ 3.4.2 ATCA monitoring
The Australia Telescope Compact Array (ATCA)444https://csiropedia.csiro.au/australia-telescope-compact-array/, located at Narrabri, New South Wales, Australia, was used to monitor AT2019aalc. The observations were conducted between 2023-06-28 and 2024-05-04, utilizing the array in the 6km (6A and 6D) and (H168) configuration. The observations were conducted at frequencies of GHz, GHz, GHz, GHz and GHz. The flux calibrator was the standard ATCA primary calibrator, 1934–638, and 1548+056 is used as phase calibrator. The observations are summarized in Table LABEL:table:atca.
The data reduction procedures were based on the Multichannel Image Reconstruction, Imaging Analysis and Display (miriad: Sault et al. 1995). The data reduction procedure was conducted in accordance with standard protocol. In order to mitigate the impact of radio frequency interference (RFI) during flux and phase calibration, the interactive flagging tool bflag was utilised. Furthermore, automated flagging routines pgflag were applied to address interference for the source. Any corrupted data identified during calibration was manually flagged.
@@
Date | Exp. time / min | Configuration | Central Freq. / GHz | Bandwidth / GHz |
28.06.2023 | 330 / 180 | 6D | 2.1 / 7.0 | 1 / 4 |
30.06.2023 | 90 / 180 | 6D | 2.1 / 7.0 | 1 / 4 |
08.07.2023 | 120 / 120 | 6D | 2.1 / 7.0 | 1 / 4 |
19.07.2023 | 210 | 6D | 2.1 | 1 |
09.08.2023 | 120 / 150 | 6D | 2.1 / 7.0 | 1 / 4 |
19.10.2023 | 180 / 180 | H168 | 2.1 / 7.0 | 1 / 4 |
09.02.2024 | 60 / 60 / 60 | 6A | 2.1 / 7.0 | 1 / 4 / 4 |
29.03.2024 | 60 / 60 / 60 | 6A | 2.1 / 7.0 / 17.0 | 1 / 4 / 4 |
@@ 3.5 Spectroscopic observations
The Low Resolution Imaging Spectrograph (LRIS: Oke et al. 1995) mounted on the Keck-I telescope at Maunakea measured the spectrum of AT2019aalc, on 2021-07-06 (PI: Kulkarni). Later, during the second optical flare, we obtained more optical spectra of AT2019aalc using the following instruments: LRIS (PI: Kulkarni), DeVeny Spectrograph mounted on the Lowell Discovery Telescope (LDT) in Arizona (PI: Hammerstein) and the Double Beam Spectrograph (DBSP, Oke & Gunn 1982) mounted on the inch telescope at Palomar Observatory, California (PIs: Kulkarni and Kasliwal). The spectra were reduced following the standard procedures using the automated reduction pipeline of the LRIS (Perley 2019), the software package PypeIt (Prochaska et al. 2020) for DeVeny, and the software packages DBSP DRT (Mandigo-Stoba et al. 2022) and PypeIt (Prochaska et al. 2020) for the DBSP. Details of these observations of AT2019aalc are summarized in Table 2 while the spectra are shown in Fig. 17.
@@
To determine the type of AGN prior the flares, we performed a stellar population synthesis fit of the pre-flare SDSS spectrum with (pPXF: Cappellari & Emsellem 2004; Cappellari 2017, 2023) and the newest GALAXEV (Bruzual & Charlot 2003) models (CB2019), assuming a Kroupa (2001) initial mass function (IMF) and an upper mass limit of , as templates (see Fig. 18). Beforehand, the spectrum was corrected for Milky Way extinction using the dust reddening maps by Schlafly & Finkbeiner (2011) and Cardelli et al. (1989) extinction curve. Furthermore, it was redshift corrected and logarithmically rebinned. We included several emission lines to the fit and assumed four kinematic components: One for all stellar templates, one for all forbidden lines, one for all allowed emission lines, and a last one for a second, broad component of all Balmer lines up to . Additionally, we included two separate dust components, each one with the extinction curve of Cardelli et al. (1989): a first one for stellar, and a second one for nebular emission. We therefore fixed the Balmer line ratio and limited doublets of the preinstalled lines to their theoretical ratio using the keywords tie_balmer and limit_doublets of pPXF. However, we only fixed the ratio for the strongest lines of the Balmer series up to to avoid possible contamination of by the [Ne iii] line. For the fit with pPXF we included additive and multiplicative polynomials up to tenth order (degree=10, mdegree=10) to model the AGN continuum. In order to correct for intrinsic dispersion caused by the spectrograph, we assume a resolving power of at and at .555https://www.sdss4.org/dr17/spectro/spectro_basics/ Linear interpolation is used to calculate values for wavelengths in between.
The newly observed spectra were fitted using the galaxy spectrum fitting tool PyQSOFit (Guo et al. 2018; Shen et al. 2019; Ren et al. 2024). The continuum of each spectrum was modeled with a combination of the following components: a power-law to represent the AGN continuum, polynomials to capture any residual smooth variations, and an Fe II template for the iron emission complex. Emission lines in the H and H complexes were modeled using a combination of broad and narrow Gaussian profiles to represent different kinematic components. The high-ionization coronal lines of interest were fitted with additional Gaussian profiles. The H complex shows a strong double-peaked feature around . This region is known as Bowen Fluorescence complex. It was fitted with a set of broad Gaussian components. We computed parameter uncertainties using MC resampling. Examples of the fitted spectra and the inferred emission line parameters are given in Fig. 19 and in Table 3, respectively.